OpenStudy (anonymous):

Write and solve the system of equations below. Let a represent adults and let c represent children. 1,200 tickets were sold to the spring carnival. Adult tickets cost $6.00 and children's tickets cost $3.50. A total of $6,450 was collected. How many tickets of each kind were sold? A. a = 800, c = 400 B. a = 900, c = 300 C. a = 1,000, c = 200 D. a = 1,100, c = 100

OpenStudy (anonymous):

Process Let x be the number of adult tickets sold and y be the number of children's tickets told. x + y = 1200, from this you know that 6x + 6y = 7200 6x + 3.5y = 6450 Subtract the two equations and you get 2.5y = 750, y = 300 x = 1200 - y = 1200 - 300 = 900

OpenStudy (anonymous):

Best answer: 900 adult tickets and 300 children's tickets were sold.

OpenStudy (kira_yamato):

1,200 tickets were sold to the spring carnival. => a + c = 1200 ---(1) Adult tickets cost $6.00 and children's tickets cost $3.50. A total of $6,450 was collected. => 6a + 3.5c = 6450 ---(2) How many tickets of each kind were sold? => Solve for a and c

OpenStudy (kira_yamato):

Answer would be 900 adult tickets, 300 children tickets.

OpenStudy (anonymous):

Hit the Best response button if I helped ;D